OpenStudy (hashuablue):

I need help with this problem: Solve for x in which the quantity plus 1 over 7 is equal to 5 over the quantity minus 1.

OpenStudy (anonymous):

\[\frac{ x+1 }{ 7 } = \frac{ 5 }{ x }-1\] or \[\frac{ x+1 }{ 7 } = \frac{ 5 }{ x-1 }\] ?

OpenStudy (hashuablue):

The second one.

OpenStudy (anonymous):

okay, so that means that (x+1)(x-1) = 35 so x^2 -1 = 35 meaning that x^2 = 35+1 since you move the 1 to the other side so x^2 = 36 so x=?

OpenStudy (hashuablue):

x=6

OpenStudy (hashuablue):

Thanks Purplerainbowcherry.

OpenStudy (anonymous):

no problem :)